# references.bib

### v1

@@ 2 Prerequisites and notation
In what follows, $\operatorname{{\mathcal{D}}}\subset\mathbb{R}^{3}$ is a smooth bounded domain with connected components $B_{1},\ldots,B_{N}$ that have connected boundaries. Their characteristic functions are denoted by $1_{B_{j}}$ and are given by
@@
$\left\{\begin{array}[]{l}\Delta u+k^{2}u=~{}~{}0\quad\text{ in }\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr u=~{}~{}\phi\quad\text{ on }\partial\operatorname{{\mathcal{D}}},\end{array}\right.$
subject to an outgoing radiation condition, has a unique solution. In that case the Dirichlet-to-Neumann operator $\mathcal{T}^{k}_{\operatorname{{\mathcal{D}}}}[\phi]$ is defined as
@@
The Dirichlet-to-Neumann operator can be expanded into a type of generalized Taylor series in the space of bounded linear operators $\mathcal{L}(H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}}),H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}}))$ for $k$ in a neighborhood of $0\in\mathbb{C}$. The following functions will be needed for the linear term of the Taylor series of $\mathcal{T}^{k}_{\operatorname{{\mathcal{D}}}}$.
Definition 2.1.
$\left\{\begin{array}[]{rll}\Delta\Phi_{j}&=~{}~{}0&\text{in }\operatorname{{\mathbb{R}^{3}}}3\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \Phi_{j}|_{\partial\operatorname{{\mathcal{D}}}}&=~{}~{}1_{\partial B_{j}}&\text{ on }\partial\operatorname{{\mathcal{D}}}.\end{array}\right.$
With these objects it is now possible to formulate the following lemma. The next results can be deduced from [feppon2024subwavelength, Proposition 3.1] and properties of the single layer potential which, e.g., can be found in [photonic].
@@
For this work, the most important part of this Lemma 2.2 will be item (iii). The formula provided for $\mathcal{T}_{0}$ and $\mathcal{T}_{1}$ will be needed for determining the subwavelength resonances asymptotically in terms of $\delta$ in Subsection 3.2 and Subsection 4.2.
In the introduction (Section 1), the fundamental theorem of subwavelength physics was presented in Theorem 1.2. The statement is that subwavelength resonances (in the linear setting) correspond to the eigenvalues of a matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ called the generalized capacitance matrix. The precise definition of the capacitance matrix is as follows [ammari2017plasmonic].
@@
Let $\operatorname{{\mathcal{D}}}\subset\mathbb{R}^{3}$ a smooth bounded domain with connected components $B_{1},\ldots,B_{N}$ that are well separated and have connected boundaries. The capacitance matrix $\operatorname{cap}(\operatorname{{\mathcal{D}}})\in\mathbb{R}^{N\times N}$ associated to $\operatorname{{\mathcal{D}}}$ is defined as
$\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}=-\int_{\partial B_{l}}\frac{\partial\Phi_{j}}{\partial\nu}\operatorname{d\sigma},$
where $\Phi_{j}$ is as in Definition 2.1. Let $\lvert B_{j}\rvert$ denote the volume of the $B_{j}\subset\operatorname{{\mathbb{R}^{3}}}3$ and ${c_{r}^{j}}$ the material parameter in $B_{j}$, for $j=1,\ldots,N$, and let $V^{vol}$ be the diagonal matrix with the diagonal entries defined as
$(V^{vol})_{j,j}=\frac{{c_{r}^{j}}^{2}}{\lvert B_{j}\rvert}.$
The generalized capacitance matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\in\mathbb{C}^{N\times N}$ associated to $\operatorname{{\mathcal{D}}}$ and the material parameter ${c_{r}}$ is defined as
$\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})=V^{vol}\operatorname{cap}(\operatorname{{\mathcal{D}}}).$
Note that if ${c_{r}^{j}}={c_{r}}$ for all $j$, then
@@
In Section 3, we will reproduce Theorem 1.2 with the techniques of Dirichlet-to-Neumann operator and a variational approach.
@@ 3 A Dirichlet-to-Neumann approach to subwavelength resonances
This section is devoted to reproducing the fundamental theorem of subwavelength physics (Theorem 1.2) with a novel technique, first used for this purpose in [feppon2024subwavelength]. The original derivation was based on layer potential techniques and the asymptotic Gohberg-Sigal theory (as it can be found for example in [photonic]). In this work, we take a different approach, which is based on Dirichlet-to-Neumann operators and variational arguments. To this end, we pose the following proposition which reformulates the system of equations (2) in terms of the Dirichlet-to-Neumann operator as in Section 2 and as a variational problem in $H^{1}(\operatorname{{\mathcal{D}}})$.
@@
An immediate consequence of Proposition 3.1 is the following lemma that characterizes variationally the resonances of the system (2).
@@
$\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}-\frac{\omega(\delta)^{2}}{{c_{r}}^{2}}\int_{\operatorname{{\mathcal{D}}}}u\overline{v}\operatorname{dvol}-\delta\int_{\partial\operatorname{{\mathcal{D}}}}\mathcal{T}^{\frac{\omega(\delta)}{{c_{0}}}}[u]\overline{v}\operatorname{d\sigma}=0,\quad\forall v\in H^{1}(\operatorname{{\mathcal{D}}}).$
@@
In the following, we will restrict our study to the subwavelength regime. That is, we will consider $\delta$ in a neighborhood of $0\in\mathbb{R}$ and $\omega$ in a neighborhood of $0\in\mathbb{C}$. We will enlarge the variational framework given in [feppon2024subwavelength] to the case of a disconnected resonator domain $\operatorname{{\mathcal{D}}}$ and characterize the subwavelength resonances in that case.
@@
is degenerate for a continuity of $(\omega,\delta)$ in a neighborhood of $0\in\mathbb{C}\times\mathbb{R}$, namely the subwavelength resonances $(\omega(\delta),\delta)$. While these characteristic points are the objects of interest, it is unfortunate that the problem is in some sense indirect. This complicates the analysis. The great advantage of the method presented in [feppon2024subwavelength] is that it translates the indirect problem of finding characteristic points to a direct problem, where one solves a well-posed variational problem. The solution $u_{\omega,\delta}$ of that problem is then used to determine whether $(\omega,\delta)$ is a resonance. Subsection 3.1 will treat the first step of developing the well-posed variational problem. In Subsection 3.2, the solution of that problem will be computed and a characterization of subwavelength resonances will be derived.
3.1 Variational formulation
@@
Define the bilinear form $a_{\omega,\delta}:H^{1}(\operatorname{{\mathcal{D}}})\times H^{1}(\operatorname{{\mathcal{D}}})\rightarrow\mathbb{C}$ associated to the smooth domain $\operatorname{{\mathcal{D}}}$ with well-separated connected components $B_{1},\ldots,B_{N}$ that have a connected boundaries,
@@
The resonances of the linear Helmholtz problem (2) will be characterized in terms of the bilinear form $a_{\omega,\delta}$ in Theorem 3.5. In the proof of Theorem 3.5, the following lemma from [feppon2024subwavelength] will be useful.
@@
There exists a neighborhood $U$ of $0\in\mathbb{R}$ and a neighborhood $V$ of $0\in\mathbb{C}$ such that for all $\delta\in U$ and for all $\omega\in V$ and for any $f\in H^{1}(\operatorname{{\mathcal{D}}})^{*}$, the dual of $H^{1}(\operatorname{{\mathcal{D}}})$, the following variational problem has a unique solution $u_{f}(\omega,\delta)$
@@
Let $\omega\in\mathbb{C}$ and $\delta\in\mathbb{R}$ belong to a neighborhood of zero. For any $f\in H^{1}(\operatorname{{\mathcal{D}}})^{*}$, the variational problem:
@@
Let $\delta_{l,j}$ denote the Kronecker delta. Integrating the above equation on $B_{l}$ gives the following system of equations:
$\begin{pmatrix}\displaystyle\int_{B_{l}}u_{f}(\omega,\delta)\operatorname{dvol}\end{pmatrix}_{l}=\sum_{j=1}^{N}\left(\delta_{l,j}-\begin{pmatrix}\int_{B_{l}}u_{1,j}(\omega,\delta)\operatorname{dvol}\end{pmatrix}_{l,j}\right)\begin{pmatrix}\int_{B_{j}}u\operatorname{dvol}\end{pmatrix}_{j},$
@@
is invertible. ∎
This theorem allows to characterize the resonances as precisely those $\omega(\delta)\in\mathbb{C}$ such that the matrix in (11) is degenerate. We have the following corollary.
Corollary 3.6.
@@
3.2 Characterization of the resonances
In this Subsection the solution of the variational problem (12) will be determined for small $\omega\in\mathbb{C}$, $\delta\in\mathbb{R}$ and then, by using Corollary 3.6, the subwavelength resonances will be computed. To this end, we consider the PDE formulation of the variational problem (12)
@@ 3.2 Characterization of the resonances
Using item (iii) in Lemma 2.2 together with Definition 2.3 and the fact from [feppon2022modal, Proposition 12] that
@@
$\mathcal{T}_{1}[1_{B_{j}}]=-\mathcal{T}^{0}_{\operatorname{{\mathcal{D}}}}\circ\mathcal{S}_{1}((\mathcal{S}^{0}_{\operatorname{{\mathcal{D}}}})^{-1}[1_{B_{j}}])=\frac{i}{4\pi}\mathcal{T}^{0}_{\operatorname{{\mathcal{D}}}}[1_{\operatorname{{\mathcal{D}}}}]\int_{\operatorname{{\mathcal{D}}}}\frac{\partial\Phi_{j}}{\partial\nu}\operatorname{dvol}=-\sum_{l=1}^{N}\frac{\partial\Phi_{l}}{\partial\nu}\sum_{l=1}^{N}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}.$
@@
$-\Delta u_{p,0}+\sum_{i=1}^{N}\left(\int_{B_{i}}u_{p,0}\operatorname{dvol}\right)1_{B_{i}}$ $=\frac{1}{{c_{r}}^{2}}u_{p-2,0}+1_{B_{j}}\delta_{p=0}\delta_{0=0},$
@@
$\left\{\begin{array}[]{rll}\Delta\tilde{u}_{0,1}^{j}&=~{}~{}0&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial\tilde{u}_{0,1}^{j}}{\partial\nu}&=~{}~{}\displaystyle\frac{1}{\lvert B_{j}\rvert}\frac{\partial\Phi_{j}}{\partial\nu}&\text{on }\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\int_{B_{l}}\tilde{u}_{0,1}^{j}\operatorname{dvol}&=~{}~{}0,&\forall~{}l\in\{1,\ldots,N\}.\end{array}\right.$
@@
$\left\{\begin{array}[]{rll}\Delta\tilde{u}_{1,1}^{j}&=~{}~{}0&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial\tilde{u}_{1,1}^{j}}{\partial\nu}&=~{}~{}\displaystyle-\frac{i}{4\pi{c_{0}}\lvert B_{j}\rvert}\sum_{l=1}^{N}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}\sum_{l=1}^{N}\frac{\partial\Phi_{l}}{\partial\nu}&\text{on }\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\int_{B_{l}}\tilde{u}_{1,1}^{j}\operatorname{dvol}&=~{}~{}0,&\forall~{}l\in\{1,\ldots,N\}.\end{array}\right.$
@@
Let $\{q_{0},v_{2},\ldots,v_{N}\}$ be a basis of eigenvectors of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ and denote by $\Pi:\mathbb{C}^{N}\rightarrow\mathbb{C}q_{0}$ the projection along $\operatorname{span}(v_{2},\ldots,v_{N})$ onto $\mathbb{C}q_{0}$. Then the subwavelength resonances $\omega(\delta)=\omega_{0}\sqrt{\delta}+\omega_{1}\delta+O(\delta^{\frac{3}{2}})$ of (2) associated to $q_{0}$ satisfies the following asymptotics:
$\omega_{0}\text{ is an eigenvector of }\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\text{ with eigenvector }q_{0},$
@@
$\left(\omega(\delta)^{2}-\delta\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})-\omega(\delta)\delta\frac{i}{4\pi{c_{0}}}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})\right)q(\delta)+O((\omega(\delta)^{2}+\delta)^{2})=0.$
@@
$\left(\omega_{0}^{2}-\delta\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\right)q_{0}=0,$
@@
It then follows that $\omega_{0}$ is an eigenvector of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ with an eigenvector $q_{0}$ and $\omega_{1}$ is given by

### v3

@@ 2 Prerequisites and notation
In what follows, $\operatorname{{\mathcal{D}}}\subset\mathbb{R}^{3}$ is a smooth bounded domain with connected components $B_{1},\ldots,B_{N}$ that have connected boundaries. Their respective volume (w.r.t. Lebesgue measure) is denoted by $\lvert B_{j}\rvert$, their characteristic functions are denoted by $1_{B_{j}}$ and are given by
@@
$\left\{\begin{array}[]{rl}\Delta u+k^{2}u&=~{}~{}0\quad\text{ in }\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr u&=~{}~{}\phi\quad\text{ on }\partial\operatorname{{\mathcal{D}}},\end{array}\right.$
subject to an outgoing radiation condition, has a unique solution. In that case, the Dirichlet-to-Neumann operator $\mathcal{T}^{k}_{\operatorname{{\mathcal{D}}}}[\phi]$ is defined as
@@
The Dirichlet-to-Neumann operator can be expanded into a type of generalized Taylor series in the space of bounded linear operators $\mathcal{L}(H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}}),H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}}))$ for $k$ in a neighborhood of $0\in\mathbb{C}$. The next results can be deduced from [12, Proposition 3.1] and properties of the single layer potential which, e.g., can be found in [5].
@@
For this work, the most important part of this Lemma 2.1 will be item (iii). The formulas provided for $\mathcal{T}_{0}$ and $\mathcal{T}_{1}$ will be needed to determine the subwavelength resonances asymptotically in terms of $\delta$ in Subsection 3.2 and Subsection 4.2.
In the introduction (Section 1), the fundamental theorem of subwavelength physics was presented in Theorem 1.2. The statement is that the leading order asymptotics of subwavelength resonances (in the linear setting) correspond to the eigenvalues of a matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ called the generalized capacitance matrix. The following functions will be needed for its definition.
Definition 2.2.
$\left\{\begin{array}[]{rll}\Delta\Phi_{j}&=~{}~{}0&\text{in }\operatorname{{\mathbb{R}^{3}}}3\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \Phi_{j}|_{\partial\operatorname{{\mathcal{D}}}}&=~{}~{}1_{\partial B_{j}}&\text{ on }\partial\operatorname{{\mathcal{D}}},\\ \Phi_{j}(x)&=\mathcal{O}(\lvert x\rvert^{-1})&\text{ as }\lvert x\rvert\rightarrow\infty.\end{array}\right.$
With these objects in hand, it is now possible to formulate a precise definition of the capacitance matrix, which is as follows [3].
@@
Let $\operatorname{{\mathcal{D}}}\subset\mathbb{R}^{3}$ be a smooth bounded domain with connected components $B_{1},\ldots,B_{N}$ that are well separated and have connected boundaries. The capacitance matrix $\operatorname{cap}(\operatorname{{\mathcal{D}}})\in\mathbb{R}^{N\times N}$ associated to $\operatorname{{\mathcal{D}}}$ is defined as
$\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}=-\int_{\partial B_{l}}\frac{\partial\Phi_{j}}{\partial\nu}\operatorname{d\sigma}.$
Let ${c_{r}^{j}}$ denote the material parameter in $B_{j}$, for $j=1,\ldots,N$, and let $V^{vol}$ and $V^{mat}$ be the diagonal matrices with the diagonal entries defined as
$(V^{vol})_{j,j}=\frac{1}{\lvert B_{j}\rvert}\quad\text{ and }\quad(V^{mat})_{j,j}=({c_{r}^{j}})^{2}.$
The generalized capacitance matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\in\mathbb{C}^{N\times N}$ associated to $\operatorname{{\mathcal{D}}}$ and the material parameters $({c_{r}^{j}})_{j}$ is defined as
$\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})=V^{mat}V^{vol}\operatorname{cap}(\operatorname{{\mathcal{D}}}).$
Note that if ${c_{r}^{j}}\equiv{c_{r}}$ for all $j$, then
@@
In Section 3, we reproduce Theorem 1.2 with the techniques of Dirichlet-to-Neumann operators and a variational approach.
@@ 3 A Dirichlet-to-Neumann approach to subwavelength resonances
This section is devoted to reproducing the fundamental theorem of subwavelength physics (Theorem 1.2) with a novel technique, first used for this purpose in [12]. The original derivation was based on layer potential techniques and the asymptotic Gohberg-Sigal theory (as can be found, for example, in [5]). In this work, we take a different approach which is based on Dirichlet-to-Neumann operators and variational arguments. To this end, we pose the following proposition which reformulates the system of equations (2) in terms of the Dirichlet-to-Neumann operator and as a variational problem in $H^{1}(\operatorname{{\mathcal{D}}})$.
@@
An immediate consequence of Proposition 3.1 is the following lemma that characterizes the resonances of the system (2) variationally.
@@
$\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}-\frac{\omega(\delta)^{2}}{{c_{r}}^{2}}\int_{\operatorname{{\mathcal{D}}}}u\overline{v}\operatorname{dvol}-\delta\int_{\partial\operatorname{{\mathcal{D}}}}\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega(\delta)}{{c_{0}}}}[u]\overline{v}\operatorname{d\sigma}=0,\quad\forall v\in H^{1}(\operatorname{{\mathcal{D}}}).$
@@
In the following, we will restrict our study to the subwavelength regime. That is, we will consider $\delta$ in a neighborhood of $0\in\mathbb{R}$ and $\omega$ in a neighborhood of $0\in\mathbb{C}$. We will extend the variational framework given in [12] to the case of a disconnected resonator domain $\operatorname{{\mathcal{D}}}$ and characterize the subwavelength resonances in that case.
@@
is degenerate for a continuity of $(\omega,\delta)$ in a neighborhood of $0\in\mathbb{C}\times\mathbb{R}$, namely the subwavelength resonances $(\omega(\delta),\delta)$. While these characteristic points are objects of interest, it is unfortunate that the problem is in some sense indirect. This complicates the analysis. The great advantage of the method presented in [12] is that it translates the indirect problem of finding characteristic points to a direct problem, where one solves a well-posed variational problem. The solution $u_{\omega,\delta}$ of that problem is then used to determine whether $(\omega,\delta)$ is a resonance. Subsection 3.1 will treat the first step of developing the well-posed variational problem. In Subsection 3.2, the solution of that problem will be computed and an asymptotic expansion of subwavelength resonances will be derived.
3.1 Variational characterization of resonances
@@
Define the bilinear form $a_{\omega,\delta}:H^{1}(\operatorname{{\mathcal{D}}})\times H^{1}(\operatorname{{\mathcal{D}}})\rightarrow\mathbb{C}$ associated to the smooth domain $\operatorname{{\mathcal{D}}}$ with well-separated connected components $B_{1},\ldots,B_{N}$ that are supposed to have connected boundaries,
@@
The resonances of the linear Helmholtz problem (2) will be characterized in terms of the bilinear form $a_{\omega,\delta}$ in Theorem 3.5. In the proof of Theorem 3.5, the following lemma from [12] will be useful.
@@
There exists a neighborhood $U$ of $0\in\mathbb{R}$ and a neighborhood $V$ of $0\in\mathbb{C}$ such that for all $\delta\in U$ and for all $\omega\in V$ and for any $f\in H^{1}(\operatorname{{\mathcal{D}}})^{*}$, the dual of $H^{1}(\operatorname{{\mathcal{D}}})$ (with respect to the $H^{1}(\operatorname{{\mathcal{D}}})$-pairing), the following variational problem has a unique solution $u_{f}(\omega,\delta)$
@@
Let $U\subset\mathbb{R}$ and $V\subset\mathbb{C}$ be open neighborhoods of 00 as in Lemma 3.4. Let $\delta\in U$ and $\omega\in V$. For any $f\in H^{1}(\operatorname{{\mathcal{D}}})^{*}$, the variational problem:
@@
Let $\delta_{l,j}$ denote the Kronecker delta. Integrating the above equation on $B_{l}$ gives the following linear system of equations:
$\begin{pmatrix}\displaystyle\int_{B_{l}}u_{f}(\omega,\delta)\operatorname{dvol}\end{pmatrix}_{l}=\sum_{j=1}^{N}\left(\delta_{l,j}-\begin{pmatrix}\int_{B_{l}}u_{1,j}(\omega,\delta)\operatorname{dvol}\end{pmatrix}_{l,j}\right)\begin{pmatrix}\int_{B_{j}}u\operatorname{dvol}\end{pmatrix}_{j},\quad l=1,\ldots,N,$
@@
is invertible.
This theorem allows us to characterize the resonances as precisely those $\omega(\delta)\in\mathbb{C}$ such that the matrix in (11) is degenerate. We have the following corollary.
Corollary 3.6 (Characterization of resonances).
@@
The proof of Theorem 3.5 says even more. In the case of a resonance, the kernel of the matrix (11) gives precisely the average in each connected component $B_{j}$ of the resonant solutions $u$.
3.2 Asymptotics of linear subwavelength resonances
In this subsection, the solution of the variational problem (12) will be determined for small $\omega\in\mathbb{C}$, $\delta\in\mathbb{R}$ and then, by using Corollary 3.6, the asymptotics of the subwavelength resonances will be computed as $\delta\rightarrow 0$. To this end, we consider the PDE formulation of the variational problem (12)
@@ 3.2 Asymptotics of linear subwavelength resonances
Using item (iii) in Lemma 2.1 together with Definition 2.3 and the fact from [11, Proposition 12] that
@@
$\mathcal{T}_{1}[1_{B_{j}}]=-\mathcal{T}^{0}_{\operatorname{{\mathcal{D}}}}\circ\mathcal{S}_{1}((\mathcal{S}^{0}_{\operatorname{{\mathcal{D}}}})^{-1}[1_{B_{j}}])=\frac{i}{4\pi}\mathcal{T}^{0}_{\operatorname{{\mathcal{D}}}}[1_{\operatorname{{\mathcal{D}}}}]\int_{\partial\operatorname{{\mathcal{D}}}}\frac{\partial\Phi_{j}}{\partial\nu}\operatorname{d\sigma}=-\sum_{l=1}^{N}\frac{\partial\Phi_{l}}{\partial\nu}\sum_{l=1}^{N}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}.$
@@
$-\Delta u_{p,0}+\sum_{i=1}^{N}\left(\int_{B_{i}}u_{p,0}\operatorname{dvol}\right)1_{B_{i}}$ $=\frac{1}{{c_{r}}^{2}}u_{p-2,0}+1_{B_{j}}\delta_{p=0},$
@@
$\left\{\begin{array}[]{rll}\displaystyle\Delta\tilde{u}_{0,1}^{j}&=~{}~{}\displaystyle-\sum_{l=1}^{N}\frac{1}{\lvert B_{j}\rvert\lvert B_{l}\rvert}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}1_{B_{l}}&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial\tilde{u}_{0,1}^{j}}{\partial\nu}&=~{}~{}\displaystyle\frac{1}{\lvert B_{j}\rvert}\frac{\partial\Phi_{j}}{\partial\nu}&\text{on }\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\int_{B_{l}}\tilde{u}_{0,1}^{j}\operatorname{dvol}&=~{}~{}0,&\forall~{}l\in\{1,\ldots,N\}.\end{array}\right.$
@@
$\left\{\begin{array}[]{rll}\Delta\tilde{u}_{1,1}^{j}&=~{}~{}\displaystyle\sum_{m=1}^{N}\frac{i}{4\pi{c_{0}}\lvert B_{j}\rvert\lvert B_{m}\rvert}(\operatorname{cap}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}}))_{m,j}1_{B_{m}}&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial\tilde{u}_{1,1}^{j}}{\partial\nu}&=~{}~{}\displaystyle-\frac{i}{4\pi{c_{0}}\lvert B_{j}\rvert}\sum_{l=1}^{N}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}\sum_{l=1}^{N}\frac{\partial\Phi_{l}}{\partial\nu}&\text{on }\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\int_{B_{l}}\tilde{u}_{1,1}^{j}\operatorname{dvol}&=~{}~{}0,&\forall~{}l\in\{1,\ldots,N\}.\end{array}\right.$
@@
Let $\{q_{0},v_{2},\ldots,v_{N}\}$ be a basis of eigenvectors of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ and denote by $\Pi:\mathbb{C}^{N}\rightarrow\mathbb{C}q_{0}$ the projection along $\operatorname{span}(v_{2},\ldots,v_{N})$ onto $\mathbb{C}q_{0}$. Then the subwavelength resonance $\omega(\delta)=\omega_{0}\sqrt{\delta}+\omega_{1}\delta+O(\delta^{\frac{3}{2}})$ of (2) associated to $q_{0}$ satisfies the following asymptotics:
$\omega_{0}\text{ is an eigenvalue of }\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\text{ with eigenvector }q_{0},$
@@
$\left(\omega(\delta)^{2}-\delta\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})-\omega(\delta)\delta\frac{i}{4\pi{c_{0}}}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})\right)q(\delta)=0.$
@@
It then follows that $\omega_{0}$ is an eigenvalue of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ with an eigenvector $q_{0}$ and $\omega_{1}$ is given by